OpenStudy (anonymous):

49x^2-25/(7x-5)=7x+5

OpenStudy (anonymous):

isit (49x^2-25)/(7x-5) = 7x + 5 or 49x^2-(25/(7x-5)) = 7x + 5

OpenStudy (king):

hey after simplifying u get 49x^2-25=49x^2-25 0=0!!

OpenStudy (mimi_x3):

Is it: \[\large \frac{49x^{2}-25}{7x-5} = 7x+5\]

OpenStudy (king):

yeah if it is what mimi posted then my answer is rite but if (7x-5) is denom. only fr 25 then answer is 8.64

OpenStudy (anonymous):

i guess it's 49x^2-(25/(7x-5)) = 7x + 5

OpenStudy (anonymous):

yea its what mimi said